# Filtering Pixel Latent Variables for Unmixing Volumetric Images

### 08 Dec 2023 v1

@@ 4.1 MNIST with simulated decay
4.2 STED-FLIM
The simulated FLIM dataset serves as a demonstration of latent unmixing in a real experimental setting, in which channel separation is challenging for low photon counts or more than two fluorescent markers. The 3D U-Net model was trained to map the input volumes of depth time-bins to a latent space of the same length, on which 4 bandpass filters equally spaced and of equal width are applied, similarly to the 3D-MNIST experiment. Our latent unmixing approach accurately predicts the pixel-wise contributions of 4 components on the test images (Fig. 5) in conditions where MLE and phasor approaches fail at correctly separating each channel (Suppl. Fig. 6 and 7). The photon counts are too low (100) for MLE to fit exponential decay curves to individual pixels; it requires more than 100 photons per pixel to accurately separate 2-color images with known lifetimes of 2 and 5 nanoseconds (Supp. sec. F). The phasor method performs better than MLE, but does not reach the capacities of our method for 4-color unmixing. It is better suited for 2- or 3- color images, which our method can also learn to unmix by adapting the number and width of filters to the target number of channels (Supp. sec. F).
| MLE | Phasors | Latent unmixing | |
| 0.9031 | 0.5046 | 0.9716 | |
| 0.0023 | 0.2601 | 0.7635 | |
| 0.0115 | 0.2577 | 0.8855 | |
| -0.0568 | 0.1327 | 0.5539 |
4.2.1 Denoising with intermediate filters
One key particularity of the latent unmixing method is that the filters applied to the latent space can be displaced to fit the needs of the task and the distribution of the latent values. For 3D-MNIST, the output values of the pixel latent space tend to form distributions with a maximum aligned with the center of the filter for each channel (Fig. 2 and 3). In the STED-FLIM experiment, the distributions tend to be skewed towards the intersection of the filters (Fig. 6). This can be interpreted as the network not being able to correctly assign a channel to a given photon, placing its contribution at the intersection of two channel-filters. We tested this hypothesis by discarding at inference the photons that are assigned to the bins surrounding the intersection of each pair of channels. This creates a fifth channel composed solely of the photons that were not confidently classified in the latent space. By removing this channel, we get an unmixed and denoised version of the input volume (Fig. 6).
4.3 Mode decomposition for multi-mode optical fibers
@@
The S-MMF dataset is used to test if the latent unmixing method can generalize to other types of distributions; here, additive sinusoidal signals with varying spatial intensities. For every single pixel, these signals are sampled under the Nyquist frequency, which is the minimal frequency of acquisition that would allow the application of Fourier-based processing methods. By applying our approach to the S-MMF volumes, the correlation from multiple pixels provide enough information to find the unknown intensity and spatial distribution of multiple coherent electromagnetic modes propagating simultaneously in a given optical fiber. The latent unmixing method was trained to map the input data to a latent space of length 60, where 6 bandpass filters of width 10 are applied to separate the space into 6 target modes (Fig. 7). Results demonstrate that the latent unmixing method can retrieve the spatial distributions of the modes (Fig. 8), which varies depending on the index profile of the fiber. To quantify the results, we compared the predicted relative intensity of each mode to the ground truth for a given sample volume. The total relative intensity is obtained by summing the pixel values over the predicted image of each mode. The predicted and ground truth fractions of intensity highly correlate 3 over all numbers of propagating modes. Interestingly, our method seems to perform better when the fiber has some intensity 4 or 5 modes than 2 or 3 modes. This goes against the trend of other methods that generally perform poorly for increased numbers of modes [24]. When there are more modes, the input volume includes more spectral variety and the 3D convolution is able to extract more information from cross-correlations gathered across the whole volume.
@@ 5 Discussion and conclusion
The latent unmixing method also offers interpretability opportunities that were briefly covered in this work. By mapping to a pixel latent space with more values than the number of channels in the ground truth (L=28 for four-channel MNIST and FLIM, L=60 for six-channel S-MMF), the network can learn to use the extra values to encode its certainty. A latent distribution that shifts from the middle of the filter can be interpreted as the network not being confident whether the photon belongs to one or the other channel. We studied this by adding narrow bandpass filters at the intersection between two filters, only at inference, for the FLIM dataset. Qualitative evaluation supports the hypothesis that photons leading to off-centered distributions seem to be associated with higher uncertainty from the network and can be used as a method of denoising images, as noise photons are more difficult to classify. This application is useful for real-life cases, like FLIM, where a background channel cannot be explicitly defined in the ground truth, as it could be for the MNIST dataset.
@@
Acknowledgments
We acknowledge the contributions of Andréanne Deschênes, Antoine Ollier and David Vincent for the scripts used to test the phasor method and maximum likelihood estimation, and Jean-Michel Bellavance for acquiring the STED images. This work was supported by the Neuronex Initiative (2014862) - National Science Foundation & Fond de Recherche du Québec - Santé, by the Natural Sciences and Engineering Research Council of Canada - Discovery Grant (DGECR-00200-2019), by the Canada Research Chair Tier 2 (CRC-0126-2019) and Canada-CIFAR AI Chair. CB and VB are both supported by NSERC PhD scholarships.
@@ A. Acquisition of microscopy images
The STED-FLIM dataset was generated from real 4-colors fluorescence microscopy images acquired with a commercial Abberior STED system. Four proteins of interest were marked with fluorescent dyes that have excitation and emission spectral signatures that allow them to be acquired sequentially. The neuronal marker MAP2 was tagged with ATTO490LS, the postynaptic marker PSD95 with STAR635, the presynaptic marker VGLUT1 with Alexa488, and the RNA binding protein FUS with Alexa594. STAR635 and Alexa594 were imaged with the STED modality (resolution 20 nm), while ATTO490LS and Alexa488 were imaged with the confocal modality (resolution 200 nm).
@@ C. Generation of the 3D-MNIST dataset
To make sure that the time dimension is exploited by the network to solve the unmixing task, a 2D version of the 3D-MNIST dataset was generated by omitting lines 23-26 and 30-33 in algo. 1; that is, by simply not creating the time dimension after creating the noisy overlapping digits. Results obtained by this 2D U-Net are shown in figure 3. The 2D U-Net can not separate the digits through the noise, demonstrating that the third dimension is useful for unmixing the overlapping digits.
@@ F. MLE and phasor method for STED-FLIM
The confusion matrices between true and predicted photon counts for each channel are shown in Supp. fig. 6(a). The MLE method tends to assign most photons to the shortest lifetime component, A.
@@
The phasor analysis method was also used to unmix the 4-color FLIM-STED dataset, and the resulting confusion matrices are shown in Supp. fig. 6(b), with example results from the FLIM-STED test set in Supp. Fig. 8 and 9.
G. Code availability
All the code used to generate the 3D-MNIST dataset, train the models, and produce the results presented here is available at https://github.com/FLClab/LatentUnmixing. All experiments were conducted using Python 3.8.10, with the library versions indicated in the repository.

### 06 Apr 2024 v2

@@ 3.1.1 Results
3.2 Fluorescence lifetime microscopy
Fluorescence lifetime microscopy (FLIM) uses the arrival time of fluorescence photons to assign them to specific fluorophores or to measure environmental changes in biological samples [6]. The fluorescence decay follows an exponential probability distribution characterized by the parameter corresponding to the fluorescence lifetime [7]. The fluorescence lifetime is a unique photophysical property of each fluorescent molecule, utilized in FLIM to differentiate between fluorophores. Despite careful selection of fluorophores with distinct fluorescence lifetime, overlapping exponential distributions pose challenges for signal unmixing. This is particularly difficult with low photon counts (less than 100 photons/pixel) or when multiple fluorophores coexist at the same location[43, 28]. For live-cell imaging, where photobleaching and sample movement should be minimized, FLIM measurements are generally associated with low photon-counts (¡25 photons per pixel[2]).
3.2.1 Related work
Different unmixing approaches have been proposed specifically for FLIM such as linear phasor analysis [13], and deep learning based strategies [42, 11, 54]. Maximum Likelihood Estimation (MLE) is a well-established curve fitting approach, that requires known components distributions when applied to multi-channel unmixing of FLIM data [4]. Phasor analysis has been proposed for the unmixing of fluorescence lifetimes, utilizing graphical representations to simplify the identification of composite lifetimes without requiring iterative fitting [13, 36]. The performance of both approaches is strongly limited by low photon counts (300) and when more than three components have to be unmixed [12, 10, 37, 13].
Recently, deep learning based approaches have been proposed for FLIM data analysis. The Fluorescence lifetime imaging network (FLI-Net) [42] is a 3D-CNN that takes as input a 3D FLIM volume (x,y,t) where each pixel is characterized by photons sampled in an exponential distribution with aim to assign a mean lifetime to each pixel rather than unmixing the input image. The fluorescence lifetime imaging method based on Generative Adversarial Network Estimation (FlimGANE)[11] uses a GAN to generate high-quality FLIM images based on photon-starved (50-200 photons per pixel) measurements. Both FLI-Net and FlimGANE models are limited to two components unmixing, since their architecture is designed for three outputs: lifetime values and and relative abundance . The lifetime of a fluorescent dye is not a fixed value and is known to vary based on the environmental conditions of the sample[44]. Training a model on a predefined set of lifetime values is impracticable, as it would unlikely capture the true measured distributions of fluorescence lifetime in a real sample.
In recent years, major developments have been made to enable multiplexed fluorescence microscopy, enabling the characterization of cellular and molecular interactions between multiple partners [16, 48, 5]. Development of analytical methods that can be applied to limited photon budget in multiplexed FLIM imaging would have an important impact in the field, allowing to increase sampling rates for live-cell imaging or to resolve nanostructures in super-resolution microscopy. Our proposed approach addresses the challenge of unmixing over up to four independent fluorophores in FLIM data, tailored for contexts where the photon count per pixel aligns with that observed in STimulated Emission Depletion (STED) super-resolution microscopy or time-lapse confocal microscopy.
3.2.2 Dataset
The STED-FLIM dataset consists of 30 real microscopy images of four different proteins tagged with fluorescent markers in fixed cortical neurons (Suppl. sec. A). The signal from each fluorescent marker was spectrally separated to obtain ground truth images with four independent channels. Prior to training, the images were split into 256 256 pixels crops using a sliding window with steps of 128 pixels. Crops were kept in the training set if the mean photon count of each of the 4 channels was at least 1 photon per pixel. This operation resulted in a set of 1806 crops for training. To augment the data, channels were randomly permuted, and up to two channels were removed. The time dimension was created during training by first uniformly sampling lifetime values , , , , and then sampling random time values for each pixels inside the dataloader from an exponential decay curve parameterized by the channels’ lifetime. This realistic dataset format could easily be adapted to fit specific parameters, number of channels, and ranges of lifetime values.
3.2.3 Training
The STED-FLIM dataset serves as a demonstration of Latent Unmixing in a real experimental setting, in which channel separation is challenging for low photon counts (¡ 50 photons per pixel), for more than two fluorescent markers, and for small lifetime difference between the emitters (¡ 1 ns). To improve the generalizability of the model, we trained a conditional U-Net inspired by the denoising diffusion model[20] that takes as input a vector in addition to the volumetric image to unmix. For targets with less than 4 channels, the empty channels are represented by . is projected to the encoding blocks by passing through two learnable fully-connected layers both with 64 nodes with a GELU activation in-between and a SiLU activation afterwards. The learned representation then goes through a third fully-connected layer with an output size corresponding to the number of channels of each encoding block (128, 256, 512 and 1024). The resulting values are then used to scale and shift the input of each encoding block. We used the implementation from https://github.com/lucidrains/denoising-diffusion-pytorch/. This model was trained on the FLIM-STED dataset with varying lifetime values and resulted in a more general model that can be applied to a wider range of lifetime pairs than an unconditional model (Suppl. sec. G). The conditional 3D U-Net model was trained to map the input volumes of depth time-bins to a latent space of the same length, on which 4 bandpass filters equally spaced and of equal width are applied, similarly to the 3D-MNIST experiment.
3.2.4 Results
Robustness to low-photon counts
(Fig. 4)
Robustness to close lifetime values
Finding pairs of fluorophores that are compatible with the proteins to be labeled and have distant lifetime values can be a challenge. The ability to correctly unmix two channels with lifetime that are close apart (¡1 ns) is therefore a valuable advantage for an unmixing method, since it opens up more possibilities of dyes that can be paired up. Our latent unmixing method shows better robustness to close lifetime than the phasor analysis method (Fig. 5), achieving a Pearson correlation of for lifetime values only 0.1 ns apart, compared to the phasor analysis method that achieves only .
4-color unmixing
Our Latent Unmixing approach accurately predicts the pixel-wise contributions of 4 components on the test images (Fig. 6) in conditions where MLE and phasor approaches fail at correctly separating each channel (Suppl. Sec. F. MLE and phasor method for STED-FLIM).
3.2.5 Application to real FLIM measurements
A key feature of the Latent Unmixing method is that it is trained on simulated data, therefore erasing the need of acquiring a FLIM dataset specifically for training. The simulated data can cover any range and combinations of lifetime values, enabling the training of a general model that can be applied to any pair of lifetime values. We applied the Latent Unmixing model trained on the simulated FLIM-STED dataset on real FLIM data (Fig. 7) with different lifetime values, photon counts, biological structures, and number of channels to separate (2, 3 and 4); all using the same general model (Suppl. sec. LABEL:). The real measurements require only very simple preprocessing steps to fit the format of the simulated training data (Suppl. sec. LABEL:).
3.3 Mode decomposition in multi-mode fibers
Multi-mode fibers are optical fibers that support the propagation of multiple simultaneous light modes [18]. Spatially and spectrally resolved (S) imaging[31] is a measurement commonly used for characterizing how light travels through multi-mode optical fibers. By measuring the spatial distribution of the intensity at the output of the fiber for a range of laser wavelengths, the modes of propagation of the light through the fiber can be characterized. The 2D image at the fiber output is not merely a linear combination of each distinct mode because the coherent modes interfere constructively or destructively with each other in function of their relative phases. In a two modes fiber, the intensity of a single pixel versus the wavelength of the laser propagating through the fiber, within a small spectral range, oscillates periodically with a sinusoidal distribution (Fig. 8). When multiple modes propagate through the fiber, the intensity oscillates as a composite of sinusoidal functions, each corresponding to the interaction between pairs of modes.
In an measurement for which the sampling frequency is above the Nyquist criterion, the measured oscillations allow to separate each mode with a Fourier transform based method [30] or by principal component analysis (PCA) [35, 39]. Other methods exist to achieve real-time mode decomposition from a single image with various algorithms [21, 3, 27], but they all require the a priori knowledge of the spatial distribution of each mode.
To our knowledge, no method for mode decomposition in multi-mode fibers allows to separate modes with unknown spatial distributions and successfully retrieve their contribution in measurements using a sub-Nyquist sampling rate.
3.3.1 Dataset
The S-MMF dataset consists in simulated images of the intensity measured at the output of optical fibers for different input wavelengths [NichoModedecompositionformulti-modeopticalfiberslson:08]. The simulated measure has three dimensions; two spatial (2D images of the light distribution at the fiber output) and one spectral (each image represents the measure obtained from a different light wavelength). The dataset was computed with the equations described in Quiquempois et al. (2021) [35]. A vector finite difference mode-solver [15] was used to find mode properties around 1550 nm for various index profiles and fiber geometries. All fibers in the dataset support up to 6 modes. The mode spatial distribution and effective index dependency on wavelength (dispersion) were taken into account. For all data samples, we considered a spectral range of 6 nm for a fiber length of 5 m. The dataset includes measurements with various combinations of up to 6 propagating modes, four different fiber geometries (step-index with varying numerical apertures and core size, GRIN, W-type, H-type and a realistic step-index geometry with a central dip), and different mode-specific intensities (). We used 1500 fiber measurements to train the network and 175 for testing.
3.3.2 Training
@@
The S-MMF dataset is used to test if the Latent Unmixing method can generalize to other types of distributions; here, additive sinusoidal signals with varying spatial intensities. For every single pixel, these signals are sampled under the Nyquist frequency, which is the minimal frequency of acquisition that would allow the application of Fourier-based processing methods. By applying our approach to the S-MMF volumes, the correlation from multiple pixels provide enough information to find the unknown intensity and spatial distribution of multiple coherent electromagnetic modes propagating simultaneously in a given optical fiber. The Latent Unmixing method was trained to map the input data to a latent space of length 60, where 6 bandpass filters of width 10 are applied to separate the space into 6 target modes (Fig. 8).
3.3.3 Results
Results demonstrate that the Latent Unmixing method can retrieve the spatial distributions of the modes (Fig. 8), which varies depending on the index profile of the fiber. To quantify the results, we compared the predicted relative intensity of each mode to the ground truth for a given sample volume. The total relative intensity is obtained by summing the pixel values over the predicted image of each mode. The predicted and ground truth fractions of intensity highly correlate 2 over all numbers of propagating modes. Interestingly, our method seems to perform better when the fiber has some intensity in 4 or 5 modes than 2 or 3 modes. This goes against the trend of other methods that generally perform poorly for increased numbers of modes [35]. When there are more modes, the input volume includes more spectral variety and the 3D convolution is able to extract more information from cross-correlations gathered across the whole volume.
@@ 4 Discussion and conclusion
The Latent Unmixing method also offers interpretability opportunities. By mapping to a pixel latent space with more values than the number of channels in the ground truth (L=28 for four-channel MNIST and FLIM, L=60 for six-channel S-MMF), the network can learn to use the extra values to encode its certainty. A latent distribution that shifts from the middle of the filter can be interpreted as the network not being confident whether the photon belongs to one or the other channel. We studied this by adding narrow bandpass filters at the intersection between two filters, only at inference, for the FLIM dataset. Qualitative evaluation supports the hypothesis that photons leading to off-centered distributions seem to be associated with higher uncertainty from the network and can be used as a method of denoising images, as noise photons are more difficult to classify. This application is useful for real-life cases, like FLIM, where a background channel cannot be explicitly defined in the ground truth, as it could be for the MNIST dataset.
@@
Acknowledgements We thank Alexy Pelletier-Rioux for the intensity images used for validation, and Albert Michaud-Gagnon and Marie Lafontaine for the implementation of the 2-species phasors unmixing method.
Declarations
4.1 Funding
Funding was provided by grants from the Natural Sciences and Engineering Research Council of Canada (NSERC) (RGPIN-06704-2019 to F.L.C. and RGPIN-2019-06706 to C.G.), Fonds de Recherche Nature et Technologie (FRQNT) Team Grant (2021-PR-284335 to F.L.C and C.G.), Sentinel North Initiative funded by Canada First Research Excellence Fund (2020-2024 Major Call for Proposal to F.L.C. and C.G.), the Canadian Institute for Health Research (CIHR) (F.L.C.), and the Neuronex Initiative (National Science Foundation 2014862, Fond de recherche du Québec - Santé 295824, ) (F.L.C.). C.G. is a CIFAR Canada AI Chair and F.L.C. is a Canada Research Chair Tier II. C.B. is supported by scholarships from NSERC, from the Fonds de Recherche Nature et Technologie (FRQNT) Quebec, from the FRQNT strategic cluster UNIQUE, and by a Leadership and Scientific Engagement Award from Université Laval. A.D is supported by … J.-M. B. is supported by …. V.B. is supported by a scholarship from NSERC.
4.2 Competing interests
The authors have no competing interests to declare that are relevant to the content of this article.
4.3 Code availability
All the code used to generate the simulated datasets, train the models, and produce the results presented here is available to download at https://anonymous.4open.science/r/LatentUnmixing-101C/. All experiments were conducted using Python 3.8.10, with the library versions indicated in the repository.
4.4
Consent for publication
Data availability
Materials availability
Author contribution
If any of the sections are not relevant to your manuscript, please include the heading and write ‘Not applicable’ for that section.
@@ A. Acquisition of microscopy images
Sample preparation
Neuronal cultures from rats were prepared following established protocols [29] in accordance with procedures approved by the animal care committee of Université Laval. Following dissociation from dissected cortexes (hippocampi for FLIM) of P0-P1 postnatal rats, cells were plated onto PDL-Laminin coated glass coverslips (12 mm) in a 24-well plate at a density of 100k cells/well (40k cells/well for FLIM). Infection with either a shFUS virus or a PLKO virus was done at 5 of 14 Days in vitro at a concentration of 5 or 10 MOI (4-color only). Fixation with 4% PFA was done at 4,7 or 11 days post infection (14 days in vitro for FLIM)
After fixation, the cells underwent three 5-minute washes in phosphate-buffered saline (PBS) supplemented with 100 mM glycine. Subsequently, the cells were permeabilized and blocked with a blocking solution composed of 0.1% Triton X-100 and 2% goat serum (GS) in PBS for 30 minutes before proceeding to immunostaining. Immunostaining involved incubation with primary antibodies (PAB) and secondary antibodies (SAB) diluted in the blocking solution at room temperature (see Supp. Table 1). Primary antibodies were allowed to incubate for 2 hours, followed by three PBS washes. Secondary antibodies were then applied for 1 hour and subsequently washed three times with PBS. Finally, coverslips were mounted onto slides using Polyvinyl alcohol mounting medium with DABCO (Sigma, cat:10981).
| Primary antibodies | ||||
| Antibody | Supplier | Catalogue no. | Dilution | Ref. |
| Mouse anti-Bassoon | Enzo | ADI-VAM-PS003 | 1 : 500 | [Dani2010, Tang2016, 52] |
| Rabbit anti-Homer1 | Synaptic Systems | 160003 | 1 : 500 | [Lagache2018, 52] |
| Mouse anti-MAP2 | Sigma-Aldrich | M4403 | 1 : 1000 | |
| Rabbit anti-FUS | Sigma-Aldrich | HPA008784 | 1 : 500 | |
| Guinea Pig anti-VGLUT1 | Sigma-Aldrich | Ab5905 | 1 : 500 | |
| Fluotag X2 Anti-PSD95 STAR 635P | Nanotag Biotechnologies | Fluotag X2 Anti-PSD95 635P | 1 : 500 | |
| Secondary Antibodies | ||||
| Antibody | Company | Catalogue no. | Dilution | |
| GAM CF594 | Sigma-Aldrich | SAB4600321 | 1 : 500 | |
| GAR STAR ORANGE | Abberior | STORANGE-1002-500UG | 1 : 250 | |
| GAM ATTO490LS | Hypermol | 2109-1MG | 1 : 250 | |
| GAGP Alexa Fluor 488 | Sigma-Aldrich | A-11073 | 1 : 250 | |
| GAR Alexa Fluor 594 | Invitrogen | A32740 | 1 : 100 | |
Imaging parameters
The STED-FLIM dataset was generated from real 4-color fluorescence microscopy images acquired with a commercial Abberior Expert Line STED microscope. Four proteins of interest were labelled with fluorescent dyes that have excitation and emission spectral signatures that allow them to be acquired simultaneously using line steps. The dendritic marker MAP2 was tagged with ATTO 490LS, the postynaptic marker PSD95 with STAR635P, the presynaptic marker VGLUT1 with Alexa Fluor 488, and the RNA binding protein FUS with Alexa Fluor 594. STAR 635P and Alexa Fluor 594 were imaged with the STED modality (resolution 20 nm), while ATTO 490LS and Alexa Fluor 488 were imaged with the confocal modality (resolution 200 nm). Further details can be found in Table 2. To acquire the real FLIM images, we equipped the microscope with a Time Tagger Ultra 8 from Swabian Instruments GmbH. The pre-synaptic protein Bassoon was tagged with CF594 and the post-synaptic protein Homer1 was tagged with STAR Orange. These two fluorescent dyes have similar excitation and emission spectra and were imaged using a single imaging channel. Further details can be found in Table 3.
| Parameter | Confocal Alexa Fluor 488 | Confocal Atto 490LS | STED Alexa Fluor 594 | STED STAR 635P |
| Excitation wavelength | 485 nm | 485 nm | 561 nm | 640 nm |
| Excitation power | ||||
| Depletion wavelength | 775 nm | 775 nm | ||
| Depletion powers | 156 mW | 156 mW | ||
| Detection wavelengths | 500-550 nm | 650-720 nm | 605-625 nm | 650-720 nm |
| Pixel dwelltime | ||||
| Line steps | 2 | 2 | 10 | 6 |
| Detection gating delay | 750 ps | 750 ps | ||
| Detection gating width | 8 ns | 8 ns | ||
| Pixel size | ||||
| Pinhole size | 1.0 A.U. | 1.0 A.U. | 1.0 A.U. | 1.0 A.U. |
| Parameter | Confocal FLIM | STED FLIM |
| Excitation wavelength | 561 nm | 561 nm |
| Excitation power | ||
| Depletion wavelength | 775 nm | |
| Depletion powers | 44, 88, 132, 176 mW | |
| Detection wavelengths | 605-625 nm | 605-625 nm |
| Pixel dwelltime | ||
| Line steps | 4 | 8 |
| Detection time bins | 250 bins of range of | 250 bins of range of |
| Pixel size | ||
| Pinhole size | 1.0 A.U. | 1.0 A.U. |
@@ C. Generation of the 3D-MNIST dataset
To make sure that the time dimension is exploited by the network to solve the unmixing task, a 2D version of the 3D-MNIST dataset was generated by omitting lines 23-26 and 30-33 in algo. 2; that is, by simply not creating the time dimension after creating the noisy overlapping digits. Results obtained by this 2D U-Net are shown in figure 3. The 2D U-Net can not separate the digits through the noise, demonstrating that the third dimension is useful for unmixing the overlapping digits.
@@ F. MLE and phasor method for STED-FLIM
The confusion matrices between true and predicted photon counts for each channel are shown in supp. fig. 6. The MLE method tends to assign most photons to the shortest lifetime component, A.
@@
The phasor analysis method was also used to unmix the 4-color FLIM-STED dataset, and the resulting confusion matrices are shown in Supp. fig. 7, with example results from the FLIM-STED test set in Supp. Fig. 8 and 9.
G. Unconditional latent unmixing for FLIM
I. Code availability
All the code used to generate the 3D-MNIST dataset, train the models, and produce the results presented here is available to download at https://anonymous.4open.science/r/LatentUnmixing-101C/. All experiments were conducted using Python 3.8.10, with the library versions indicated in the repository.